Cannon shell impact position, as measured along the line of fire from the target point, can be described by a normal random variable X. It is found that 15.15% of shells fall 11.2 m or farther from the target in a direction toward the cannon, while 5.05% fall farther than 95.6 m beyond the target. what are the mean and standard deviation of distribution of X?


1
Expert's answer
2021-02-04T05:33:07-0500

Let "X=" the cannon shell impact position, as measured along the line of fire from the target point: "X\sim N(\mu, \sigma)."

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0,1)"


"P(X<-11.2)=P(Z<\dfrac{-11.2-\mu}{\sigma})=0.1515"

"\dfrac{-11.2-\mu}{\sigma}\approx-1.03002"



"P(X>95.6)=1-P(X\leq95.6)="

"=1-P(Z\leq\dfrac{95.6-\mu}{\sigma})=0.0505"

"P(Z\leq\dfrac{95.6-\mu}{\sigma})=0.9495"

"\dfrac{95.6-\mu}{\sigma}\approx1.64002"

"95.6-1.64002\sigma=-11.2+1.03002\sigma"

"2.67004\sigma=106.8"

"\sigma=40"

"\mu=95.6-1.64002(40)=30"

"\mu=30\ m,\sigma=40\ m"